OpenStudy (anonymous):

NEED HELP! Making a scale model of earth were 1 mm is 20 km, how tall (in mm) would these measurements be: 22.5 km 2867 km 2266 km 1216 km

jimthompson5910 (jim_thompson5910):

22.5 km: (1 mm)/(20 km) = (x mm)/(22.5 km) 1/20 = x/22.5 1*22.5 = 20x 22.5 = 20x 20x = 22.5 x = 22.5/20 x = 1.125 So 22.5 km is represented by 1.125 mm on the scale model

jimthompson5910 (jim_thompson5910):

You do the same with each measurement

OpenStudy (anonymous):

Oh Okay thanks! thats what i was doing but its huge! haha sorry! anyway thanks so much!

jimthompson5910 (jim_thompson5910):

np